Experimental Study of General Circulations with the Parameters Relevant to Earth: General Circulations on an Aqua-Planet

∗Yoshi-Yuki Hayashi^{1}, Shin-ichi Takehiro^{2}, Kensuke Nakajima^{2}, Masahiro Hosaka^{3}, and Masaki Ishiwatari^{4} (^{∗}Contact Person)

1Grad. School of Math. Sci., Univ. of Tokyo, Komaba 3-8-1, Meguro-ku, Tokyo, 153-8914, Japan

and Grad. School of Sci., Hokkaido Univ., Kita 10 Nishi 8, Sappro, 060-0810, Japan

2Faculty of Science, Kyushu Univ. Hakozaki 6-10-1, Fukuoka, 812-81, Japan

3Meteofrolgical Research Institute, Japan Met. Agency, Nagamine 1-1, Tsukuba, 305-0052, Japan 4Grad. School of Env. Earth Science, Hokkaido Univ., Kita 10 Nishi 5, Sappro, 060-0810, Japan

Abstract

“Aqua planet” experiments are performed in order to investigate the effects of an equatorial warm sea surface temperature (SST) area on the tropical large-scale precipitation patterns. The warm SST area is placed at the equator of an hypothetical aqua planet whose SST distribution is zonally uniform and symmetric about the equator.

The calculated tropical precipitation distributions are characterized by the appearance of an east-west asymmetry; precipitation decreases to the west of the warm SST area, while it increases in the longitudinally wide area to the east. In the western region to the warm SST area, an increase of the stability due to a temperature rise in the middle layer is observed and correspondingly downward flow is recognized. In the eastern region, although a temperature rise appears in the middle layer, the stability decreases due to the increase of water vapor in the lower layer caused by the meridional moisture convergence. Experiments in which the wind velocity used in the evaluation of the surface evaporation is fixed are performed to show that WISHE (Wind Induced Surface Heat Exchange) mechanism is not a principal cause for the generation of the east-west asymmetry of the precipitation pattern.

The appearance of the decrease of precipitation to the wet of the warm SST area may explain the occurrence of drought over the Indonesian region during the El Nino events.

Keywords: Aqua planet, Warm SST anomaly, Equatorial precipitation, WISHE, ENSO

Background

The grand purpose of our study is to reveal the dynamical structures which underlie the general circulations of the earth and planetary atmospheres. It is aimed to construct a theoretical framework which is useful in describing their circulation characteristics, as tried by Golitsyn (1970). Through these activities, our interest is to recognize the atmosphere of the earth (or the climate of the earth) as one of the possible realizations in the physical parameter space observed in the solar system.

One of our procedures in revealing the possible underlying dynamics is to gather and classify atmospheric circulation patterns which might be observed under various values of planetary “external” conditions such as orbital parameters (amount and variation of the incoming solar flux), radius and rotation rate of the planet, radiation property of the atmosphere, and surface boundary setups. The sampling of the possible atmospheric circulations is to a certain extent possible by numerical experiments by the use of the super computer powers.

Our search in circulation patterns is now performed for the following three major activities:

1. to reveal the possible circulations which might be realized with the values of solar flux and orbital parameters which are related to Mars, Earth and Venus,

2. to reveal the possible circulations which might be realized with the earth’s condition, but with strongly simplified surface and/or physical processes,

3. to reveal the possible circulations which might be realized as convection of a spherical shell in general.

1 E-mail address: shosuke@gfd-dennou.org

These are being studied continuously from the preceding years by our research group.

The present study belongs to the second activity listed above. The hypothetical earth utilized in this study is so called aqua-planet, where all of the earth’s surface is assumed to be covered by the ocean (Hayashi and Sumi, 1986; Numaguti and Hayashi, 1991a, b; Numaguti, 1993). The aim is to answer the basic problem of the climate, that is, “where and how does it rain?”. In the aqua-planet experiments performed so far, the focus has been placed on especially in searching for the idealistic precipitation distribution of the tropics. The experiment performed here is also in this category, however, the new feature is to find out the deformation of the precipitation distribution to a warm SST (sea surface temperature) anomaly placed in the tropics. The configuration is supposed to extract the effect of the warm SST region in the real Western Pacific on the general circulation patterns.

We will show here the precipitation patterns over the entire tropical region caused by a local warm SST area placed at the equator of an aqua planet. On the warm SST area, an intense precipitation activity, so called, convection center, will emerge. However, we will consider the precipitation patterns all over the tropical region affected by the existence of the convection center; we will not focus on the convection center itself. The SST distribution outside the warm SST area is zonally uniform and symmetric about the equator to minimize the in situ effect on the precipitation properties.

If we can neglect the various feedback mechanisms in the moist atmosphere, the precipitation patterns outside the convection center will be easily predicted. The downward motion should be observed in the neighborhood of the convection center, which should suppress precipitation there. However, we know that it is incomplete to predict the precipitation patterns without considering the water vapor supply and its transport. In the following results obtained by the numerical experiments, it is actually shown that precipitation decreases to the west of the convection center, but it increases to the east (Hosaka et al., 1998).

Model

The model used in this study is AGCM5 of GFD Dennou Club edition, which is the threedimensional primitive system on a sphere (Swamp Project, 1998). The model is basically the same as that developed by and used in Numaguti (1993). The dynamical part is represented by the pseudo spectral method with the triangular truncation at wavenumber 42 (T42) and by the sigma vertical coordinate with 16 vertical levels.

The model contains the following simplified physical processes. The employed cumulus parameterization is the Kuo scheme or the adjustment scheme. The condensed water is removed from the system immediately after condensation and hence there are no clouds. The absorption and scattering of the solar radiation are not included. The longwave absorption is represented by four spectral bands, three of which are of water vapor and one is of dry air. The scattering of the longwave radiation is not included. The absorption coefficients of the longwave radiation are chosen so that the cooling profile of the atmosphere roughly resembles the observed one. The vertical turbulent mixing is represented by the level II scheme of Mellor and Yamada. The surface fluxes of momentum, heat and water vapor are evaluated by the bulk formula.

In this report, we will describe only the results with adjustment scheme. It has been confirmed that the obtained tendency of the deformation of precipitation pattern is independent of the choice of cloud parameterization.

Figure 1: The distribution of SST used in experiment A4. The local warm SST area is shaded. The figure in the right hand side shows the basic SST distribution (solid line) and the distribution at the longitude 180^{∘}, where the value of SST anomaly is the maximum (dashed line).

Experimental Design

The surface is all covered by the ocean (aquaplanet) and the value of SST is fixed. The basic SST distribution is the same as that used by Numaguti (1993). It is symmetric about the equator and zonally uniform. The value of SST at the equator is 302K and remains constant in the equatorial region between the latitudes of ± 7^{∘}. The shape of the local warm SST area is rectangular with the longitudinal extent of 40^{∘} and the latitudinal extent of 20^{∘}. The peak value of SST anomaly is 2, 4, 6 or 8K, and is located at the center of the warm SST area which is placed at the equator. Accordingly, the experiment without SST anomaly is referred to as A0 and the experiments with SST anomaly are referred to as A2, A4, A6, and A8 respectively. Fig.1 shows the SST distributions of A0 and A4.

For the experiment A0, the model is integrated from an isothermal rest state and the data from day 700 to 2200 are used for analyses. For the warm SST area experiments (A2 - A4), the model is integrated from the data at day 700 of A0 respectively, and the data from day 150 to 900 are used for analyses. The model atmosphere is recognized to reach a quasi-equilibrium state after 150 days of integration.

Results: Precipitation Patterns

Fig.2 is the time mean horizontal distribution of the precipitation difference for experiment A4 from the zonal mean distribution of experiment A0.

Figure 2: Time mean horizontal distribution of precipitation difference for experiment A4 from the zonal mean distribution of experiment A0. Unit is W/m^{2}. Contour interval is 20W/m^{2}.

The notable feature in Fig.2 is the deformation of ITCZ. There appears an east-west asymmetry in the precipitation distribution at the equator. To the west of the warm SST area, the amount of precipitation decreases in the region with the longitudinal extent of about 70^{∘}. To the east of the warm SST area, on the other hand, it increases in the region with the longitudinal extent of about 220^{∘}. The precipitation anomaly has its minimum of −100 W/m^{2}in the western region at the longitude of 150^{∘} and its maximum of +60 W/m^{2}in the eastern region at the longitudes of 220^{∘} and 280^{∘}. In the regions along ±10^{∘} latitudes, the amount of precipitation decreases regardless of the eastern or western sides. Precipitation increases only at the longitudes where the warm SST is placed.

As for experiments A2, A6 and A8 we show only the longitudinal distributions of precipitation at the equator (Fig.3). The east-west asymmetry in the precipitation pattern observed in experiment A4 is also seen in these experiments; the amount of precipitation decreases in the western region while it increases in the broad eastern region. As the warm SST anomaly increases, the amount of precipitation at the convection center increases. However, the increase of precipitation in the eastern region seems to be independent of the magnitude of the SST anomaly. Even a small SST anomaly of experiment A2 causes a precipitation increase of the same magnitude as that of experiment A4. On the other hand, the drying in the western region is intensified as the warm SST anomaly increases. The western region of experiment A2 is not so dry as that of experiment A4. The drying in experiments A6 and A8 is intensified compared to that of experiment A4. Note that in experiments A6 and A8 the decrease of precipitation occur even in the eastern vicinity of the warm SST area.

Figure 3: Longitudinal distributions of precipitation at the equator in the warm SST area experiments. Thin solid line, thick solid line, thin dashed line and thick dashed line indicate experiments A2, A4, A6 and A8, respectively. The zonal mean precipitation value of experiment A0 is subtracted. Unit is W/m^{2}.

The East-West Asymmetry

A necessary condition for occurrence of moist convection is that the value of θ_{e} in the lower level atmosphere is higher than the minimum of θ_{e}^{∗} (usually located in the middle troposphere). According to the numerical results, θ_{e}^{∗} of the warm SST experimnts increases in the middle troposphere (σ ∼ 0.55) both to the east and to the west of the warm SST area, and the amount of increase is the larger in the eastern region. Since the increase of θ_{e}^{∗} reflects the warming in the middle troposphere, its overall increase is qualitatively consistent with the intuitive view in which the atmosphere around a convection center is stabilized by the middle tropospheric warming. However, the fact that the eastern region is more stable (warmer) than the western region implies that precipitation should be more strongly suppressed to the east than to the west. The important result is that a significant east-west asymmetry can be observed in the change of θ_{e} in the lower levels. This zonally asymmetric response is consistent with that of precipitation.

The increase of low level θ_{e} to the east of the warm SST area consists mainly of the increase of humidity; the increase of θ_{e} due to the temperature change is about 0.2K while that due to the moisture change is about 1K. The increase of the low level humidity to the east of the warm SST area should be accounted for either by the increase of evaporation from the ocean or by that of horizontal moisture convergence.

(a)

(b)

Figure 4: Horizontal distributions of (a) evaporation and (b) P − E of experiment A4. Unit is W/m^{2}. The zonal mean values of experiment A0 are subtracted, respectively.

Now, we will evaluate the relative importance of the two candidates, WISHE and CISK, by examining the response of evaporation and that of the difference between precipitation and evaporation (indicated by P −E below) which reflects the horizontal moisture convergence. Fig.4(a) shows the difference in evaporation (E) in the tropics between experiments A4 and A0, and Fig.4(b) is that for P − E. Around ±7^{∘} latitudes to the east of the warm SST area, a slight increase of evaporation is found. However, against the previous expectation on WISHE, the equatorial response to the east of the warm SST area is negative. On the other hand, P − E increases at the equator to the east of the warm SST area. As P − E equals a horizontal moisture convergence in the equilibrium state, this indicate the existence of significant horizontal moisture convergence at the equator to the east of the warm SST area. It is evident that the change of evaporation does not play an important role in moistening at the equator.

(a)

(b)

Figure 5: Longitudinal distributions of precipitation (solid line) and evaporation (dashed line) of experiment A4 (a) at the equator and (b) averaged in the meridional area between ±30^{∘} latitudes. Unit is W/m^{2}. The zonal mean values of experiment A0 are subtracted, respectively.

In order to show the responses more clearly, the changes of evaporation and precipitation at the equator are plotted in Fig.5(a). Again, it is evident that precipitation is enhanced to the east of the warm SST area and suppressed to the west. The change in evaporation, on the other hand, is very small compared to that of precipitation except for over the warm SST area.

Fig.5(b) shows the longitudinal distributions of the changes of precipitation and evaporation averaged meridionally between ±30^{∘} latitudes. There is no east-west asymmetry in the evaporation response, just as at the equator. WISHE does not operate even as the meridional average over the whole tropical latitudes.

By comparing Fig.5(a) and Fig.5(b), we can recognize that the increase of precipitation in the eastern region at the equator (Fig.5a) is compensated by the decrease of that in the subtropics, resulting in little change in the meridionally averaged precipitation (Fig.5b). This suggests that the equatorial precipitation enhancement to the east of the warm SST area is accounted for by the change of meridional moisture transport. The enhancement of precipitation at the equator by the change of the meridional moisture flux to the east of the warm SST area can be recognized also in the other experiments with different amplitudes of the warm SST anomaly.

In order to evaluate the effect of WISHE more completely, experiment A4VC is conducted in which WISHE is excluded artificially. In experiment A4VC, the surface wind speed used in this evaporation bulk formula is held constant in time in the tropical region between the latitudes of ± 30^{∘}. The fixed value of constant wind speed is obtained from the temporal and zonal mean of experiment A0 except for the warm SST area where the temporal mean value of experiment A4 is used. The initial condition of experiment A4VC is the data at day 2200 of experiment A0 and is integrated for 900 days and the data of the last 500 days is used for the following analyses (not shown here). The result showed that the precipitation at the equator to the east of the warm SST area increases significantly from that of experiment A0. The response in P −E is quite similar to that of experiment A4, which means that the precipitation response is strongly connected to the horizontal moisture convergence. WISHE plays at most secondary role in the generation of the zonally asymmetric response of precipitation.

Concluding Remarks

By a series of aqua planet experiments with a localized warm SST area with various intensities, following conclusions are obtained. The response to a warm SST area placed at the equator is that precipitation is enhanced in an extensive region to the east of the warm SST area while drying occurs to the west. The increase of precipitation in the eastern region results from the increase of meridional moisture convergence associated with the frictional inflow converging into the low pressure area existing to the east of the warm SST area. The decrease of precipitation in the western region is associated with downward flow and an increase of stability due to a temperature rise in the middle layer. WISHE does not play an important role in the generation of the east-west asymmetry of the precipitation response.

By comparing the results of the two series of experiments with the different cumulus parameterization schemes (the adjustment scheme and the Kuo scheme), the structure of the precipitation response is found to be insensitive to the choice of cumulus parameterization scheme. The east-west asymmetric response of precipitation obtained in the present study is robust in this sense. The characteristic that an intense dry region appears to the west of an active convection center is found in the real tropical atmosphere and the GCM experiments. During the ENSO warm events, strong droughts are observed to the west of the warm SST anomaly areas in the equatorial Pacific (Ropelewski and Halpert 1987). In the 82–83 ENSO warm event, the decrease of cumulus activity is clearly shown by the OLR observation (e.g., Arkin et al. 1983); it is reported that forest fires occurred even in the tropical rain forest of Borneo (Tsuyuki, 1988) ^{1} . In the GCM SST anomaly experiments, dry areas also appear to the west of the anomalous regions (e.g., Keshavamurty, 1982; Geisler et al., 1985; Wu and Liu, 1992). As for the climatic precipitation distribution, there appear small rainfall regions to the west of the convection centers (Spencer, 1993). For instance, the heavy convection activity over the Maritime Continent does not extend westward far into the Indian Ocean along the equator. The reason for the small rainfall is usually considered as a result of low SST values in the corresponding region. However, even if the values of SST to the eastern and western sides are the same, the results of present research show that a dry region extends to the west of the convection center. The characteristics exemplified in our numerical calculations contribute at least partly to the formation of precipitation patterns to the west of the real convection centers.

Acknowledgment

We wish to thank all of the staff members of CGER for providing us with this powerful computing facility. Our software environment was provided from the resources of GFD Dennou Club (http://www.gfd-dennou.org/).

1 The forest fires of 97 event are recently reported to be even more severe.

References

Arkin, P.A., Kopman, J.D. and Reynolds, R.W. (1983) 1982-1983 El Nino / Southern Oscillation event quick look atlas. NOAA / National Weather Service, National Meteorological Center, Climate Analysis Center, Washington D. C. 20323.

Geisler, J. E., Blackmon, M. L., Bates G. T. and Munoz, S. (1985) Sensitivity of January climate response to the magnitude and position of equatorial Pacific sea surface temperature anomalies. J. Atmos. Sci., 42, 1037-1049.

Golitsyn, G.S. (1970) A similarity approach to the general circulation of planetary atmospheres. ICARUS, 13, 1-24.

Hayashi, Y.-Y. and Sumi, A. (1986) The 30-40 day oscillations simulated in an “aqua planet” model. J. Met. Soc. Japan, 64, 451-467.

Hosaka, M., Ishiwatari, M., Takehiro, S., Nakajima, K. and Hayashi, Y.-Y. (1998) Tropical precipitation patterns in response to a local warm SST area placed at the equator of an aqua planet. J. Met. Soc. Japan, 76, 289-305.

Keshavamurty, R. N. (1982) Response of the atmosphere to sea surface temperature anomalies over the equatorial Pacific and the teleconnections of the southern oscillation. J. Atmos. Sci., 39, 1241-1259.

Numaguti, A. and Hayashi, Y.-Y. (1991) Behaviors of cumulus activity and the structures of circulations in an “aqua planet” model. Part I: The structure of super clusters. J. Met. Soc. Japan, 69, 541-561.

Numaguti, A. and Hayashi, Y.-Y. (1991) Behaviors of cumulus activity and the structures of circulations in an “aqua planet” model. Part II: Eastward moving planetary scale structure and the intertropical convergence zone. J. Met. Soc. Japan, 69, 563-579.

Numaguti, A. (1993) Dynamics and energy balance of the Hadley circulation and the tropical precipitation zones: Significance of the distribution of evaporation. J. Atmos. Sci., 50, 1874-1887.

Ropelewski, C. F. and Halpert, M. S. (1987) Global and Regional Scale Precipitation Patterns Associated with the El Nino/Southern Oscillation. Mon. Wea. Rev., 115, 1606-1626.

Spencer, R. W. (1993) Global oceanic precipitation from the MSU during 1979–91 and comparisons to other climatologies. J. Climate, 6, 1301-1326.

Swamp Project (1998) AGCM5. http://www.gfd-

dennou.org/arch/agcm5, GFD Dennou Club. Tsuyuki, S. (1988) エルニーニョ現象と世界の天候. 昭

和 63 年度長期予報課研修テキスト— 熱帯の対流活動と日本の天候 —. 長期予報課昭和 63 年 9 月気象庁予報部, 10-16. (in Japanese).

Wu, G.-X. and Liu, H.-Z. (1992) Atmospheric precipitation in response to equatorial and tropical sea surface temperature anomalies. J. Atmos. Sci., 49, 2236-2255.